OpenStudy (anonymous):

Definition for cot theta? 1 cot theta=tan^-1 theta 2 cot theta=sin theta /cos theta 3 cot theta=cos theta/sin theta 4 cot theta=1 theta/cos theta 5 cot theta=1/sin theta

OpenStudy (studygurl14):

\(\Large \cot(\theta)=\frac{1}{\tan(\theta)}\) If, \(\large \tan(\theta)=\Large\frac{\sin(\theta)}{\cos(\theta)}\), what does \(\large \cot(\theta)\) equal? Hint: plug in the value of tan(theta) into the the first equation and simplify.

OpenStudy (anonymous):

Number 3?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Thanks

OpenStudy (studygurl14):

Yep. Nice job!